# DESI EMISSION LINE GALAXIES: CLUSTERING DEPENDENCE ON STELLAR MASS AND [OII] LUMINOSITY

### v2

@@ 2.3 Aperture Correction
Every spectrum measured by DESI undergoes a wavelength-dependent flux correction using previously identified stars as calibration standards (Guy et al., 2023). However, because ELGs are spatially extended, the fraction of light captured by DESI’s 1.5′′-diameter fiber differs from that of the standard stars, meaning the amplitude of the calibration vector is not accurate. An accurate flux calibration for galaxies requires corrections for the seeing at the time of observation and the surface brightness profile.
We apply a customized aperture correction to the [OII] line flux for ELGs used in this study. To do so, we must first back out the point-source assumption in the original calibration. We assume an average seeing of 1′′ FWHM, for which the DESI fiber will capture $\sim$79of a standard star’s flux and yield an original scale factor of 1.266. After factoring this correction out, we apply our customized aperture correction using the DESI Legacy Survey’s $r$-band photometry (Dey et al., 2019), with the central wavelength corresponding to $\sim$3,100 Å in the rest-frame of a galaxy at $z=1$. We multiply the (now uncorrected) flux by the ratio of total $r$-band flux to the predicted amount within the DESI fiber in 1′′ FWHM seeing. The median of this ratio across our samples is $\sim$1.59. Functionally, our total procedure is
@@ 3 Clustering Measurements
We first quantify the clustering of galaxies with the two-point correlation function, which describes the excess probability of finding a galaxy pair with transverse separation $r_{p}$ and line-of-sight separation $r_{\pi}$. The 2PCF is measured with respect to a random distribution that shares the same footprint as the galaxy sample. We use the Landy-Szalay estimator to make the measurement (Landy & Szalay, 1993),
@@
Here, $DD$ is the fraction of all data-data pairs in a separation bin centered on $\left(r_{p},\>r_{\pi}\right)$. $DR$ and $RR$ are similarly defined but for data-random and random-random pairs. In the transverse direction, evenly spaced logarithmic bins with width of $\Delta\log\left[r_{p}/\left(h^{-1}\rm{Mpc}\right)\right]\sim 0.21$ were used. Along the line-of-sight, evenly spaced linear bins with $\Delta r_{\pi}=4$ $h^{-1}$pc were used to a maximum value of $r_{\pi,\>\rm max}=60$ $h^{-1}$pc.
@@
PIP weighting was done by following the procedure outlined in Lasker et al. (2025). 128 alternative realizations of the DESI targeting algorithm are performed. From these, each galaxy is assigned a bitwise completeness weight—a series of 128 bits where each bit indicates whether the galaxy was observed (bit value of 1) or not observed (bit value of 0) in a realization. The PIP weight of a galaxy pair is one over the fraction of realizations in which both galaxies were observed (Bianchi & Percival, 2017),
@@
where $w_{m}^{(b)}$ is the bitwise weight of galaxy $m$, $N_{\rm real}$ is the total number of realizations, and popcount($w_{m}^{(b)}\>\&\>w_{n}^{(b)}$) counts the number of realizations in which both galaxies were observed. Here, $N_{\rm real}=129$ to account for the 128 alternative realizations and the survey observation itself. An offset of one is added in the denominator for the same reason.
@@
We then integrated $\xi(r_{p},\>r_{\pi})$ along the line-of-sight to calculate the projected 2PCF, which minimizes effects from redshift space distortions (e.g., Hamilton 1998),
@@
For each galaxy sample, the covariance matrix was estimated via 160 jackknife samples, each omitting one eighth of one survey rosette. Elements of the matrices were calculated following Zehavi et al. (2005),
@@
Clustering and number density measurements for the low, intermediate, and high redshift samples are shown in Figures 4-6, respectively.
Figure 4: Measurements and fits for the nine $1.00.8<z<$ 1.0 samples defined in Figure 3 and Table 1. The top-left, top-right, and bottom-left panels show the projected 2PCFs of the low, intermediate, and high stellar mass samples, respectively. Filled circles mark the measurements. Solid curves and shaded regions mark the median and $1\sigma$ bounds of the posterior model distribution, respectively. For illustrative purposes, measurements are staggered in $r_{p}$. Vertical offsets of $-1$ dex and $+1$ dex are applied to the lowest and highest [OII] luminosity samples in each panel, respectively. The sub-panel below each main panel shows the residual of the model relative to the measurement, normalized by the measurement uncertainty. The shaded region represents the ±$\pm 1\sigma$ bounds. Bottom-right: Measurements of the number densities are shown as filled circles, with 10% uncertainties assumed. Open circles and error bars mark the median and $1\sigma$ bounds of the posterior model distribution, respectively. For illustrative purposes, horizontal offsets are added between the measurement and fit of a given sample and between samples.
Figure 5: Same as Figure 4 but for the $1.21.0<z<$ 1.2 intermediate-redshift samples.
Figure 6: Same as Figure 4 but for the $1.61.2<z<$ 1.6 high-redshift samples.
4 Modeling the [OII] Luminosity- and Stellar Mass-Dependent Clustering
@@ 4.1 Occupation Model for Galaxies
Following the CCMD framework in Xu et al. (2018) and Clontz et al. (2022), we describe the conditional $\log L_{\rm[OII]}-\log M_{*}$ distribution of galaxies occupying halos of mass $M_{h}$; centrals and satellites are each given a distinct model. The mean number of central and satellite galaxies occupying halos of a given halo mass are denoted by $\langle N_{\rm cen}(M_{h})\rangle$ and $\langle N_{\rm sat}(M_{h})\rangle$, respectively. For ease of reading, we define the logarithmic stellar mass as $x\equiv\log M_{*}$ and the logarithmic [OII] luminosity as $y\equiv\log L_{\text{[OII]}}$. Across the central and satellite models, we have 15 free parameters to describe the conditional $\log L_{\rm[OII]}-\log M_{*}$ distribution of ELGs. A summary of the model parameters is given in Table 2. We provide a detailed description in what follows.
@@ 4.1.1 Distribution of Central Galaxies
Here, $A_{c}$ is the fraction of halos with mass $M_{h}$ that are occupied by a central ELG, $\mu_{x}$ and $\mu_{y}$ are the respective median values of $\log M_{*}$ and $\log L_{\text{[OII]}}$, $\sigma_{x}$ and $\sigma_{y}$ are the respective standard deviations, and $\rho$ is the correlation coefficient between $\log M_{*}$ and $\log L_{\text{[OII]}}$. In this work, we assume $\log M_{*}$ and $\log L_{\text{[OII]}}$ are uncorrelated at fixed halo mass, meaning $\rho=0$.
The parameters in the above 2D Gaussian equations depend on halo mass following similar descriptions in Xu et al. (2018) and Clontz et al. (2022). The amplitude of the Gaussian varies as a power law,
@@ 4.1.2 Distribution of Satellite Galaxies
We also describe the $\log L_{\rm[OII]}-\log M_{*}$ distribution of satellites as a function of halo mass. First considering only stellar mass, the conditional (logarithmic) stellar mass function of satellite galaxies is described by a modified Schechter function in similar vein to Yang et al. (2008),
@@
$=\phi_{s}10^{(\alpha_{s}+1)(x-x_{s})}\exp\left[-10^{2(x-x_{s})}\right],$ (15)
where $\phi_{s}$ is a normalization, $\alpha_{s}$ is the slope at low stellar mass, $M_{*,s}$ is the characteristic mass that marks the transition between power-law and exponential behavior, and $x_{s}\equiv\log{M_{*,s}}$ ). Our model has the flexibility to describe a gap between the satellite characteristic stellar mass ($x_{s}$) and the median central stellar mass ($\mu_{x}$). In this work, we fix that gap to zero, making $x_{s}=\mu_{x}$.
To incorporate [OII] luminosity, $\log{}L_{\text{[OII]}}$ is described as a Gaussian at fixed $M_{*}$, with median value $\mu_{y,\>\rm sat}$ and width $\sigma_{y,\>\rm sat}$. Thus the $\log L_{\rm[OII]}-\log M_{*}$ distribution of satellites becomes
@@
We parameterize the halo mass dependence of $\phi_{s}$ in a similar way to Equation (10),
@@
Motivated by the star formation main sequence (Sparre et al., 2015), we parameterize $\mu_{y,\>\rm sat}$ to be linearly dependent on $x$,
@@
$\log M_{h,p}$ Pivot halo mass for the ELG fraction-halo mass relation 11
$\log A_{c,p}$ ELG fraction at the pivot halo mass $\left[-2,0\right]$
$\gamma_{A}$ Power-law index for the ELG fraction-halo mass relation $\left[-1,1\right]$
[…]
$\alpha_{x}$ Power-law index for the stellar-halo mass relation $\left[0.25,0.31\right]$
$\log\sigma_{x}$ Logarithmic scatter in $\log M_{*}$ $\left[-1,0\right]$
$\log M_{ty}$ Transition halo mass for the [OII] luminosity-halo mass relation $\left[5,14\right]$
[…]
$\log\sigma_{y}$ Logarithmic scatter in $\log L_{\rm[OII]}$ $\left[\log 0.005,0\right]$
$\rho$ $\log L_{\rm[OII]}-\log M_{*}$ correlation coefficient 0
@@
[…]
$\gamma_{\phi}$ Power-law index for the normalization-halo mass relation 1
$\alpha_{s}$ Slope at low stellar mass $\left[-5,0\right]$
$x_{s}$ Characteristic stellar mass of power law-to-exponential transition $\mu_{x}$
$x_{p}$ Pivot stellar mass for the [OII] luminosity-stellar mass relation 10
[…]
$\log\sigma_{y,\>\rm sat}$ Logarithmic scatter in $\log L_{\rm[OII]}$ $\left[\log\sigma_{y},0\right]$
@@
$\log M_{tx}-\log h-\log\left(1-\alpha_{x}\right)$ Peak location of the stellar-halo mass ratio $\left[11.477,12.301\right]$
$\mu_{x,t}-\log M_{tx}+\log h$ Peak value of the stellar-halo mass ratio $\left[\log 0.005,\log 0.05\right]$
$\mu_{y,\>\rm sat}$ at $\log M_{h}=12$ Median $\log L_{\rm[OII]}$ of satellites $\left[41,\infty\right)$
$\frac{\text{d}\langle N_{\rm sat}(M_{h})\rangle}{\text{d}x}$ at $\log M_{*}=10$ Conditional stellar mass function of satellites $\left[10^{-4},\infty\right)$
@@ 4.2 Calculation of Observables and Derived Quantities
The satellite fraction, $f_{\rm sat}$, of each sample is calculated as
@@
For a given redshift, the model simultaneously fits 108 measurements (nine $\log L_{\rm[OII]}-\log M_{*}$ samples times 12 measurements per sample) with 15 free parameters, resulting in 93 degrees of freedom. The $\chi^{2}$ of the global fit is calculated as
@@
Here, $\boldsymbol{w_{p,\>\rm data}}$ ($\boldsymbol{w_{p,\>\rm model}}$) is the projected 2PCF data (model) vector, $\boldsymbol{n_{\rm data}}$ ($\boldsymbol{n_{\rm model}}$) is the number density data (model) vector, and $\bf{C}_{\boldsymbol{w_{p}}}$ and $\bf{C}_{\boldsymbol{n}}$ are the respective covariance matrices. We assume 10% uncertainty in the measured number density of each sample. To explore the model parameter space, we employ a Markov chain Monte Carlo (MCMC) method with a likelihood function proportional to $\exp(-\chi^{2}/2)$.
@@ 5 Results and Discussion
At fixed redshift, the measured number densities are roughly equal by construction, as shown in the bottom-right panels of Figures 4-6. The projected 2PCF measurements (other panels of Figures 4-6) have several notable features. First, they lack strong trends with respect to stellar mass, [OII] luminosity, or redshift. This absence is likely a consequence of the measurement scatter in stellar mass ($\sim$0.5 dex as shown in Appendix A) and a lack of correlation between [OII] luminosity and stellar mass (Favole et al., 2020). Second, clustering measurements on the smallest scales ($r_{p}\lesssim 0.6$ $h^{-1}$pc) have large uncertainties. This reflects the small survey volume and low number densities of our samples, resulting in a dearth of close pairs. Finally, at fixed redshift and on scales above $r_{p}\sim 10$ $h^{-1}$pc, the projected 2PCFs show various degrees of spurious signal, deviating from the expected nearly constant galaxy bias. For example, the measurements for low and intermediate stellar mass samples are not parallel to each other. This may suggest contamination caused by poorly-understood imaging systematics.

### v3

@@ 2.3 Aperture Correction
Every spectrum measured by DESI undergoes a wavelength-dependent flux correction using previously identified stars as calibration standards (Guy et al., 2023). However, because ELGs are spatially extended, the fraction of light captured by DESI’s 1.5′′ diameter fiber differs from that of the standard stars, meaning the amplitude of the calibration vector is not accurate. An accurate flux calibration for galaxies requires corrections for the seeing at the time of observation and the surface-brightness profile.
We apply a customized aperture correction to the [OII] line flux for ELGs used in this study. To do so, we must first back out the point-source assumption in the original calibration. We assume an average seeing of 1′′ FWHM during spectroscopic observation, for which the DESI fiber will capture $\sim$79% of a standard star’s flux and yield an original scale factor of 1.266. After factoring this correction out, we apply our customized aperture correction using the DESI Legacy Survey’s $r$-band photometry (Dey et al., 2019), with the central wavelength corresponding to $\sim$3100 Å in the rest frame of a galaxy at $z=1$. We multiply the (now uncorrected) flux by the ratio of total $r$-band flux to the predicted amount within the DESI fiber in 1′′ FWHM seeing. The median of this ratio across our samples is $\sim$1.59. Functionally, our total procedure is
@@ 3 Clustering Measurements
We first quantify the clustering of galaxies with the 2PCF, which describes the excess probability of finding a galaxy pair with transverse separation $r_{p}$ and line-of-sight separation $r_{\pi}$. The 2PCF is measured with respect to a random distribution that shares the same footprint as the galaxy sample. We use the Landy-Szalay estimator to make the measurement (Landy & Szalay, 1993):
@@
Here, $DD$ is the fraction of all data-data pairs in a separation bin centered on $\left(r_{p},\>r_{\pi}\right)$. $DR$ and $RR$ are similarly defined but for data-random and random-random pairs. In the transverse direction, evenly spaced logarithmic bins with width of $\Delta\log\left[r_{p}/\left(h^{-1}\rm{Mpc}\right)\right]\sim 0.21$ were used. Along the line of sight, evenly spaced linear bins with $\Delta r_{\pi}=4$ $h^{-1}$Mpc were used up to a maximum value of $r_{\pi,\>\rm max}=60$ $h^{-1}$Mpc.
@@
PIP weighting was done by following the procedure outlined in Lasker et al. (2025). One-hundred and twenty-eight alternative realizations of the fiber assignment algorithm are performed, in which each target’s subpriority (used to split priority ties between targets of the same class) is assigned a new random number. From these realizations, each galaxy is assigned a bitwise completeness weight—a series of 128 bits where each bit indicates whether the galaxy was observed (bit value of 1) or not observed (bit value of 0) in a realization. The PIP weight of a galaxy pair is one over the fraction of realizations in which both galaxies were observed (Bianchi & Percival, 2017):
@@
[…]
@@
We then integrated $\xi(r_{p},\>r_{\pi})$ along the line of sight to calculate the projected 2PCF, which minimizes effects from redshift space distortions (e.g., Hamilton 1998):
@@
For each galaxy sample, the covariance matrix was estimated via 160 jackknife samples, each omitting one-eighth of one survey rosette. Elements of the matrices were calculated following Zehavi et al. (2005):
@@
Clustering and number density measurements for the low-, intermediate-, and high-redshift samples are shown in Figures 4-6, respectively.
Figure 4: Measurements and fits for the nine $0.8<z<1.0$ samples defined in Figure 3 and Table 1. The top-left, top-right, and bottom-left panels show the projected 2PCFs of the low-, intermediate-, and high-stellar-mass samples, respectively. Filled circles mark the measurements. Solid curves and shaded regions mark the median and $1\sigma$ bounds of the posterior model distribution, respectively. For illustrative purposes, measurements are staggered in $r_{p}$. Vertical offsets of 1-1 dex and +1+1 dex are applied to the lowest and highest [OII] luminosity samples in each panel, respectively. The subpanel below each main panel shows the residual of the model relative to the measurement, normalized by the measurement uncertainty. The shaded region represents the $\pm 1\sigma$ bounds. Bottom right: measurements of the number densities are shown as filled circles, with 10% uncertainties assumed. Open circles and error bars mark the median and $1\sigma$ bounds of the posterior model distribution, respectively. For illustrative purposes, horizontal offsets are added between the measurement and fit of a given sample and between samples.
Figure 5: Same as Figure 4 but for the $1.0<z<1.2$ intermediate-redshift samples.
Figure 6: Same as Figure 4 but for the $1.2<z<1.6$ high-redshift samples.
4 Modeling the [OII] Luminosity- and Stellar-mass-dependent Clustering
@@ 4.1 Occupation Model for Galaxies
Following the CCMD framework in Xu et al. (2018) and Clontz et al. (2022), we describe the conditional $\log L_{\rm[OII]}-\log M_{*}$ distribution of galaxies occupying halos of mass $M_{h}$; centrals and satellites are each given a distinct model. The mean number of central and satellite galaxies occupying halos of a given halo mass are denoted by $\langle N_{\rm cen}(M_{h})\rangle$ and $\langle N_{\rm sat}(M_{h})\rangle$, respectively. For ease of reading, we define the logarithmic stellar mass as $x\equiv\log\left[M_{*}/M_{\odot}\right]$ and the logarithmic [OII] luminosity as $y\equiv\log\left[L_{\rm{[OII]}}\,/\,\left(\rm{erg\;\,s}^{-1}\right)\right]$. Across the central and satellite models, we have 15 free parameters to describe the conditional $\log L_{\rm[OII]}-\log M_{*}$ distribution of ELGs. A summary of the model parameters is given in Table 2. We provide a detailed description in what follows.
@@ 4.1.1 Distribution of Central Galaxies
Here, $A_{c}$ is the fraction of halos with mass $M_{h}$ that are occupied by a central ELG, $\mu_{x}$ and $\mu_{y}$ are the respective median values of $\log M_{*}$ and $\log L_{\text{[OII]}}$, $\sigma_{x}$ and $\sigma_{y}$ are the respective standard deviations (which describe the combination of intrinsic scatter and measurement uncertainty), and $\rho$ is the correlation coefficient between $\log M_{*}$ and $\log L_{\text{[OII]}}$. In this work, we assume $\log M_{*}$ and $\log L_{\text{[OII]}}$ are uncorrelated at fixed halo mass, meaning $\rho=0$.
The parameters in the above 2D Gaussian equations depend on halo mass following similar descriptions in Xu et al. (2018) and Clontz et al. (2022). The amplitude of the Gaussian varies as a power law:
@@ 4.1.2 Distribution of Satellite Galaxies
We also describe the $\log L_{\rm[OII]}-\log M_{*}$ distribution of satellites as a function of halo mass. First considering only stellar mass, the conditional (logarithmic) stellar mass function of satellite galaxies is described by a modified Schechter function in a similar vein to Yang et al. (2008):
@@
$\phi_{s}10^{(\alpha_{s}+1)(x-x_{s})}\exp\left[-10^{2(x-x_{s})}\right],$ (15)
[…]
@@
We parameterize the halo mass dependence of $\phi_{s}$ in a similar way to Equation (10):
@@
Motivated by the star formation main sequence (Sparre et al., 2015), we parameterize $\mu_{y,\>\rm sat}$ to be linearly dependent on $x$:
@@
$\log M_{h,p}$ Pivot halo mass for the ELG fraction-halo mass relation. 11
$\log A_{c,p}$ ELG fraction at the pivot halo mass. $\left[-2,0\right]$
$\gamma_{A}$ Power-law index for the ELG fraction-halo mass relation. $\left[-1,1\right]$
$\log M_{tx}$ Transition halo mass for stellar-halo mass relation. $\left(-\infty,\infty\right)$
$\mu_{xt}$ Median $\log M_{*}$ at the transition halo mass. $\left(-\infty,\infty\right)$
$\alpha_{x}$ Power-law index for the stellar-halo mass relation. $\left[0.25,0.31\right]$
$\log\sigma_{x}$ Logarithmic scatter in $\log M_{*}$. $\left[-1,0\right]$
$\log M_{ty}$ Transition halo mass for the [OII] luminosity-halo mass relation. $\left[5,14\right]$
$\mu_{yt}$ Median $\log L_{\rm[OII]}$ at the transition halo mass. $\left(-\infty,\infty\right)$
$\alpha_{y}$ Power-law index for the [OII] luminosity-halo mass relation. $\left(-\infty,\infty\right)$
$\log\sigma_{y}$ Logarithmic scatter in $\log L_{\rm[OII]}$. $\left[\log 0.005,0\right]$
$\rho$ $\log L_{\rm[OII]}-\log M_{*}$ correlation coefficient. 0
@@
$\log\phi_{s,p}$ Normalization at the pivot halo mass. $\left(-\infty,\infty\right)$
$\gamma_{\phi}$ Power-law index for the normalization-halo mass relation. 1
$\alpha_{s}$ Slope at low stellar mass. $\left[-5,0\right]$
$x_{s}$ Characteristic stellar mass of power law-to-exponential transition. $\mu_{x}$
$x_{p}$ Pivot stellar mass for the [OII] luminosity-stellar mass relation. 10
$\mu_{yp,\>\rm sat}$ Median $\log L_{\rm[OII]}$ at the pivot stellar mass. $\left(-\infty,\infty\right)$
$\gamma_{ys}$ Power-law index for the [OII] luminosity-stellar mass relation. $\left(-\infty,\infty\right)$
$\log\sigma_{y,\>\rm sat}$ Logarithmic scatter in $\log L_{\rm[OII]}$. $\left[\log\sigma_{y},0\right]$
@@
$\log M_{tx}-\log h-\log\left(1-\alpha_{x}\right)$ Peak location of the stellar-halo mass ratio. $\left[11.477,12.301\right]$
$\mu_{x,t}-\log M_{tx}+\log h$ Peak value of the stellar-halo mass ratio. $\left[\log 0.005,\log 0.05\right]$
$\mu_{y,\>\rm sat}$ at $\log M_{h}=12$ Median $\log L_{\rm[OII]}$ of satellites. $\left[41,\infty\right)$
$\frac{\text{d}\langle N_{\rm sat}(M_{h})\rangle}{\text{d}x}$ at $\log M_{*}=10$ Conditional stellar mass function of satellites. $\left[10^{-4},\infty\right)$
@@ 4.2 Calculation of Observables and Derived Quantities
The fraction of ELGs that are satellites (known as the satellite fraction) in each sample, $f_{\rm sat}$, is calculated as
@@
For a given redshift, the model simultaneously fits 108 measurements (the 2PCF in 11 $r_{p}$ bins, plus the number density, for each of nine $\log L_{\rm[OII]}-\log M_{*}$ samples) with 15 free parameters, resulting in 93 degrees of freedom. The $\chi^{2}$ of the global fit is calculated as
@@
Here, $\boldsymbol{w_{p,\>\rm data}}$ ($\boldsymbol{w_{p,\>\rm model}}$) is the projected 2PCF data (model) vector, $\boldsymbol{n_{\rm data}}$ ($\boldsymbol{n_{\rm model}}$) is the number density data (model) vector, and $\bf{C}_{\boldsymbol{w_{p}}}$ and $\bf{C}_{\boldsymbol{n}}$ are the respective covariance matrices. We assume 10% uncertainty in the measured number density of each sample. To explore the model parameter space, we employ a Markov Chain Monte Carlo method with a likelihood function proportional to $\exp(-\chi^{2}/2)$.
@@ 5 Results and Discussion
At fixed redshift, the measured number densities are roughly equal by construction, as shown in the bottom-right panels of Figures 4-6. The projected 2PCF measurements (other panels of Figures 4-6) have several notable features. First, they lack strong trends with respect to stellar mass, [OII] luminosity, or redshift. This absence is likely a consequence of the measurement scatter in stellar mass ($\sim$0.5 dex as shown in Appendix A) and a lack of correlation between [OII] luminosity and stellar mass (Favole et al., 2020). Second, clustering measurements on the smallest scales ($r_{p}\lesssim 0.6$ $h^{-1}$Mpc) have large uncertainties. This reflects the small survey volume and low number densities of our samples, resulting in a dearth of close pairs. Finally, at fixed redshift and on scales above $r_{p}\sim 10$ $h^{-1}$Mpc, the projected 2PCFs show various degrees of spurious signal, deviating from the expected nearly constant galaxy bias. For example, the measurements for low- and intermediate-stellar-mass samples are not parallel to each other. This may suggest contamination caused by poorly understood imaging systematics.